OpenStudy (anonymous):

what must you add to make it a complete square? x^2+8x=11?

OpenStudy (anonymous):

you got this yes? half of 8 is 4, 4 squared is 16 done

OpenStudy (anonymous):

+5

OpenStudy (anonymous):

16

OpenStudy (anonymous):

add 16 on both sides to make it a perfect square on the left hand side. (x+4)^2=27

OpenStudy (anonymous):

I just don't understand how to get these answers.... I wasn't taught this...

OpenStudy (anonymous):

to get a perfect square see the single degree x term and see if u can seperate into equal half if so den multiply those term u will get a answer and den see wat number u can add to get that value hope ull get it

OpenStudy (anonymous):

yah you just need to know the formula (a+b)^2=a^2+2ab+b^2 you should get your equation in the form of a^2+2ab+b^2 from the question x^2+8x=11 x^2+8x-11=0 simplify like this x^2+(2)(4)(x)-11=0 from that a=x and b=4 so -11 should be replaced by b^2 to make it perfect square so to get b^2=16 add 27 . then it becomes x^2+8x+16=0